OpenStudy (anonymous):

write an explicit formula for the geometric sequence 25, 5, 1, 1/5

OpenStudy (anonymous):

Do you see a relation between each number and the one next to it?

OpenStudy (anonymous):

yes, they are divided by 5

OpenStudy (anonymous):

there is your formula

OpenStudy (anonymous):

xn+1= xn /5 what it's saying is that the next number is 1/5 the last number

OpenStudy (anonymous):

so that would be \[A _{n-1}=\frac{ A _{n} }{ 5 }\]

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

you're welcome, always willing to help